# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ Abstract
The latest developments in computer hardware, sensor technologies, and artificial intelligence can make virtual reality (VR) and virtual spaces an important part of human everyday life. Eye tracking offers not only a hands-free way of interaction but also the possibility of a deeper understanding of human visual attention and cognitive processes in VR. Despite these possibilities, eye-tracking data also reveals users’ privacy-sensitive attributes when combined with the information about the presented stimulus. To address all these possibilities and potential privacy issues, in this survey, we first cover major works in eye tracking, VR, and privacy areas between 2012 and 2022. While eye tracking in the VR part covers the complete pipeline of eye-tracking methodology from pupil detection and gaze estimation to offline use of the data and analyses, as for privacy and security, we focus on eye-based authentication as well as computational methods to preserve the privacy of individuals and their eye-tracking data in VR. Later, considering all of these, we draw three main directions for the research community by focusing on privacy challenges. In summary, this survey provides an extensive literature review of the utmost possibilities with eye tracking in VR and the privacy implications of those possibilities.
Index Terms:
virtual reality, eye tracking, privacy, security, survey, literature review.
@@ I Introduction
Over the last decade, virtual and augmented reality (VR/AR) communities have benefited from computer hardware, graphics, and imaging science developments. To date, some modern head-mounted displays (HMDs) have already become available for reasonable prices for everyday usage. Furthermore, eye-tracking sensors have become either directly integrated into these HMDs (e.g., HTC Vive Pro Eye [1] or are available as low-cost add-ons (e.g., Pupil Labs’ eye-tracker [2]). Since eye movements are related to human cognition, visual attention, and perception, the information on where users look at a certain point in time can be helpful to users in various ways, such as providing them with adaptive support during their tasks in VR, increasing their engagement, or improving the usability of the VR applications. At the same time, in the area of VR and HMDs, there is also a lot of room for improvement, such as enhancing display resolutions, increasing realism, or preventing cybersickness. One can potentially utilize gaze information to mitigate some of these technical issues.
It was foreseen by NVIDIA in 2017 that VR was approximately 20 years away from resolutions similar to human eyes [3]. With such technical developments and a wider range of usage, future HMDs may become mobile devices like today’s mobile phones, tablets, or smartwatches. However, as these devices are situated on the heads of users and have close proximity to the eyes of the users, apart from user-supporting functionalities, they might be perceived as privacy-invasive by users since a lot of sensitive attributes about users such as genders, sexual preferences, and personal identities, can be extracted using eye movements [4]. From a privacy perspective, this information should be protected, or the user should have control over its release. At the same time, a pleasant virtual experience should be available to all users. In addition to possibilities with eye movements, eye information can also be utilized for authentication purposes with high accuracy, especially with iris textures, as they are analogous to fingerprints. Therefore, regarding privacy, there is an essential utility-privacy trade-off that practitioners should consider, along with user preferences and concerns.
As VR and privacy communities have been working to tackle these issues, and with the availability of HMDs to wider communities, especially in recent years, the number of works in the intersection of VR, eye tracking, and privacy has been increasing. However, until now, none of the previous works have focused on comprehensive coverage of eye tracking in VR and its privacy implications by considering the leading venues in the field. In this work, (1) we first cover the research between 2012 and 2022 by first discussing how eye movements are extracted and the possibilities of using eye movements in immersive VR. Then, (2) we discuss security and privacy implications, including eye-based authentication and privacy-preserving eye tracking. Lastly, (3) we draw and discuss three different directions for future research, especially considering the privacy aspects of immersive eye-tracked VR.
@@ I-A Paper’s Aim, Structure, and a Navigational Guide
This paper aims to provide researchers and practitioners with a detailed overview of the different aspects of eye tracking in VR. We start with details on existing hardware, datasets, and algorithms, the latter being more particular about eye-region segmentation and gaze estimation. We then dig deeper into eye-based interaction techniques in VR, which often inherently utilize estimated eye regions and gaze. These methods help with utilizing eye movements in real-time in VR to track the users’ attention or for interaction purposes, and we aim to equip the reader with information on the existing hardware, datasets, and algorithms. Furthermore, the broad overview of the works about human cognition, visual attention, and perception in VR, either in real-time or offline, aims to teach the reader how human eye movements and physiology are related. Security and privacy aspects discussed in the paper aim to provide methodological details on the relationship between eye movements and user authentication, and how to protect the users’ privacy when the focus is not on authentication. Lastly, future directions bring a privacy-aware take on eye movement-based methods in VR. Overall, the main goal of this paper is to provide a thorough guide to navigating the landscape of eye tracking in VR, covering its technical underpinnings, practical uses, and ethical implications while also suggesting directions for further exploration, especially for privacy protection.
@@
As we comprehensively survey research conducted in eye tracking, VR, and privacy domains, and due to the interdisciplinary and disparate nature of each domain, we present the structure of this paper and provide a navigational guideline on how readers with different backgrounds can benefit from it. We first discuss existing works in the literature in Section II. We then discuss the most important academic venues where research in these domains is published and our methodology for filtering relevant papers from each venue to include them in our survey in Section III. These sections are appropriate for all readers, especially those who want to understand existing surveys and how this work differs from them.
@@
Section IV discusses a range of topics, from computer vision-based approaches to track the eyes and further process eye region data to eye-based human-computer interaction and understanding human visual attention, cognition, and perception. Section IV-A will benefit most readers interested in computer vision and machine learning on how to track eyes and make eye-tracking solutions available for others. Section IV-B focuses on human-computer interaction by utilizing eye movements for different purposes, such as real-time interaction in VR, foveated rendering, and how to deal with technical issues to improve immersive experiences. This section is more suitable for researchers and practitioners who work on the technical aspects of human-computer interaction and eye tracking. Furthermore, Section IV-C combines eye tracking and VR from a human visual attention, cognition, and perception point of view and is more suitable for researchers with cognitive and experimental psychology backgrounds or who are interested in conducting human factors research.
In Section V, considering all the works and insights provided in Section IV, we provide works that revolve around the privacy and security implications of eye-tracked VR, including authentication possibilities using eye movement data, why privacy-preserving methods are essential for this domain, and how current literature has computationally addressed the privacy issues. This section is relevant both for privacy and security researchers and for human-computer interaction researchers interested in eye-tracked VR’s privacy and security aspects. Lastly, even though we provide future directions and discussions mainly on privacy implications of eye-tracked VR in Section VI, we note that this section is very relevant for all the researchers and practitioners whose work includes VR and eye tracking to a certain extent as researchers and practitioners should handle privacy issues carefully regardless of the research area or application domain. In Section VII, we conclude our paper. In summary, after reading this paper either wholly or partly, we foresee that it will be easier for a novice individual to be knowledgeable about the state-of-the-art research in the field. At the same time, our overview will serve as a compact reference to the experienced researcher.
@@ II Related work
Several works analyze relevant research for eye tracking, VR/AR, privacy, and security. Despite this, none of the works systematically and comprehensively review eye-tracking in VR setups by covering the major venues, the privacy issues that such works could lead to for the end users, and how to mitigate such risks. Previously, Lappi [5] analyzed eye tracking in the wild by discussing the advantages, disadvantages, and technical terms to achieve valid and high-quality results from eye-tracking experiments. Duchowski [6] discussed gaze applications and gaze-based interactions in graphical systems by focusing on eye movement analytics, foveated rendering, and visual attention in a SIGGRAPH course. In another work, Plopsky et al. [7] covered gaze-based interaction and eye tracking in head-worn extended reality (XR). Furthermore, Silva et al. [8] discussed the foundations of eye-tracking support for visual analytics systems, applications, and challenges by five critical themes, including privacy protection as an essential and challenging issue to be solved, as eye-tracking data can even reveal disorders with high accuracy.
Ens et al. [9] analyzed the challenges in immersive analytics by considering virtual, mixed, and augmented realities and drew future directions. Furthermore, Li and Barmaki [10] provided recent research on eye tracking for VR and AR by identifying visual attention, collaborative learning, emotion detection, and privacy as the main points. Sector- and domain-specific reviews, such as Pastel et al.’s work [11], focused on the eye-tracking systems for immersive VR and their possible transfer to the sports sector. The authors stressed the possibility of improving usability, interaction, and image presentation when VR integrates eye-tracking systems. With a similar focus, Heilmann and Witte [12] provided a literature review on the differences between stimulus presentations and motor responses in eye-tracking studies, particularly emphasizing sports. The authors indicated the advantages of VR in terms of experimental control and repeatability.
As this work focuses on VR setups, especially for mixed and augmented reality solutions, we refer the reader to the previous works in this area [13, 14, 15]. Related to VR, Clay et al. [16] examined eye-tracking research in VR along with a pilot user study by focusing on experimental setups, common problems of virtual environments such as vergence-accommodation conflict [17], eye-tracking calibration, and visual region of interests. However, their study was small scale and not very comprehensive. Recently, Adhanom et al. [18] provided a broad overview of eye-tracking applications in VR and their challenges. Like Silva et al.’s work [8], the authors identified privacy and security as an essential discussion point. However, their review was not systematic and covered privacy as an application with a handful of works in the field rather than comprehensively reviewing the previous works and discussing the implications. Moreno-Arjonilla et al. [19] conducted a similar review on eye tracking in VR and focused on eye-tracking processing pipeline details such as calibration and gaze estimation, with its applications in different domains such as medicine and neuroscience. However, the authors did not provide any insights on privacy. Gressel et al. [20] provided a brief overview of the privacy aspects of eye tracking and recommendations for privacy-aware eye tracking. Furthermore, overviews on considerations for high-quality data collection [21] and measurement of data quality of HMD-based eye tracking based on accuracy and precision [22] also exist in the literature, which is important for privacy because when privacy-preserving solutions are in place, a privacy-utility trade-off is often regarded and assessed.
@@
To improve data analytics, quality, and evaluations, while visual scanning patterns and eye-tracking data consist of useful information about users, such information also treats users’ privacy. Liebling and Preibusch [4] discussed that the end-users may not understand the privacy loss and argued that privacy protection should be in place for gaze and pupillometry. Kröger et al. [23] pointed out the attributes that can be inferred using gaze data similar to Liebling and Preibusch’s work [4] and discussed the privacy implications and possible societal impacts. Lebeck et al. [24] considered multi-user AR’s security and privacy aspects. The authors found that some users were concerned about the powerful abilities of the eye-tracking enabled interfaces and behavioral tracking, such as an AR HMD understanding that the user is being attracted to someone due to being unable to stop looking at them, which might also apply to VR. This statement partly overlaps with the findings of more recent work by Steil et al. [25] that users agree to share their eye-tracking data if the co-owner of the data is a governmental health agency or the purpose is research. Furthermore, Katsini et al. [26] provided an extensive survey on gaze-based authentication, and like other previous works [18, 8], encouraging further research for privacy-preserving eye tracking.
@@
Despite several recent attempts, none of the previous works have focused on systematically analyzing the research that studies eye-tracking data analytics in VR, considering the whole processing pipeline of eye-tracking methods. We consider methods from pupil detection and gaze estimation to human visual attention and cognition for eye-tracking data processing and mining. We further cover works that study privacy-preserving manipulations of eye movements in VR setups and provide further novel research directions that should be pursued to help VR technology be used in everyday life. In addition, various papers in the literature identified privacy as one of the application fields. However, considering the number of sensitive attributes that eye-tracking data can reveal, we take privacy as one of the main pillars of our paper. Therefore, our survey, focusing on these aspects, provides a comprehensive and systematic overview for the research community.
@@ III Methodology
As a starting point, we analyzed the relevant surveys and the papers that study eye tracking, VR, AR, human-computer interaction, and privacy-preserving eye tracking [5, 8, 9, 7, 13, 4, 23, 26]. Following, we considered the papers between 2012-2022 since the VR devices and eye trackers have been becoming prevalent in the daily life lately. We used Google Scholar to query papers from 35 renowned venues including ACM CHI, IEEE VR, ACM ETRA, IEEE ISMAR, ACM IUI, ACM UIST, ACM ICMI, ACM MobileHCI, ACM MM, ACM VRST, IEEE AIVR, AAAI Conference on Artificial Intelligence, IEEE CVPR, IEEE ICCV, ECCV, NeurIPS/NIPS, ACM SUI, IEEE VIS, PoPETS/PETS, USENIX Security, USENIX SOUPS, NDSS, IEEE Symposium on S&P, ACM CCS, ACM Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), Journal of Eye Movement Research (JEMR), IEEE Transactions on Visualization and Computer Graphics (TVCG), ACM Transactions on Graphics (ToG), IEEE S&P Magazine, Journal of Privacy and Confidentiality, IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Dependable and Secure Computing, Computers & Security, ACM Transactions on Privacy and Security, and ACM Transactions on Cyber-Physical Systems. For the conference proceedings, we only considered the main conference proceedings, leaving companion and workshop proceedings aside, apart from the works that present concrete results or that are highly related to eye tracking and VR. Proceedings of the ACM ETRA in 2022 were published in the Proceedings of the ACM on Human-Computer Interaction and Proceedings of the ACM on Computer Graphics and Interactive Techniques, and we also included these papers in our search results under the ETRA category. We cross-checked the query results from the DBLP.
To filter the relevant papers, we used (“eye tracking” AND “virtual reality”) as the main query, where each phrase could appear anywhere in the papers. In addition, we used the following sub-queries to further track privacy- and biometrics-related papers: (“eye tracking” AND “virtual reality” AND “privacy”) and (“eye tracking” AND “virtual reality” AND “biometrics”). We also applied forward and backward tracking by reviewing the citations and references of the queried papers, respectively. The main query yielded 1364 results, whereas two sub-queries led to 215 and 96 results, respectively. Figure 1 shows the resulting distributions of the main query according to venues. If a venue above does not appear in Figure 1, it means that we did not find any paper at those venues with our main query.
Figure 1: Query responses according to venues for the query (’eye tracking’ AND ’virtual reality’) between 2012-2022.
@@ IV Eye-tracked Virtual Reality
Eye tracking can be considered one of the key technologies, especially for VR setups, as it is possible to track eyes more accurately within HMDs compared to remote setups due to the reduced distance between eye trackers and eyes in HMDs. In addition, as eyes do not move completely voluntarily, one may likely get more objective measurements for user behaviors compared to self-reported data, such as questionnaires. While one can utilize this rich source of information in various ways, there are different technical details. Firstly, to use this data, eye regions and gaze directions are extracted. These tasks are mostly done using computer vision and machine learning techniques. Later, estimated gaze directions are especially utilized for interaction purposes, for instance, to support users in various ways, such as by predicting user intents and providing them with context-sensitive aid or purely technical ways, such as foveated rendering. While such cases should work in real-time, collected data from different experimental setups can also be used to understand humans and how they behave in various application domains, such as education or medicine. These application domains indeed show that VR can also be considered a research tool to create scenarios that are almost impossible to realize in the real world due to privacy and safety issues or simply because of practical infeasibility.
Considering these, we organize this section as follows. We first discuss works that utilize computer vision and machine learning to estimate gaze and detect eye regions, as these are the initial steps of the eye-tracking data processing. A significant portion of the recent research on these utilized deep learning-based approaches. We further include the eye movement event detection methods in VR, which differ from the conventional settings. Then, we provide a detailed overview of eye-based interaction and its different applications. In most of these works, efficiency, usability, and real-time working capability were considered essential points in eye-based human-computer interaction. As the last step, we cluster the works that conduct offline data analyses, particularly to understand human visual attention, cognition, and perception, and we provide an overarching view of these works. While the offline understanding of human visual attention, cognition, and perception does not have a direct and real-time effect on the interaction experience, understanding those aspects helps for an optimal virtual space and interaction design. We provide the structure of this section in Figure 2. In addition, we provide the visual representation of eye-tracking data capturing, data processing, data types, and utility in a holistic manner in Figure 3.
Figure 2: Organization of Section IV.
Figure 3: Visual representation of eye-tracked VR, considering data capturing, processing, data types, and utility. The visualization does not exhaustively include all the possibilities.

### v3

@@ Abstract
The latest developments in computer hardware, sensor technologies, and artificial intelligence can make virtual reality (VR) and virtual spaces an important part of human everyday life. Eye tracking offers not only a hands-free way of interaction but also the possibility of a deeper understanding of human visual attention and cognitive processes in VR. Despite these possibilities, eye-tracking data also reveals users’ privacy-sensitive attributes when combined with the information about the presented stimulus. To address all, this survey first covers major works in eye tracking, VR, and privacy areas between 2012 and 2022. While eye tracking in VR part covers the computational eye tracking pipeline from pupil detection and gaze estimation to offline data analysis, for privacy and security, we focus on eye-based authentication as well as computational methods to preserve the privacy of individuals and their eye-tracking data in VR. Later, we outline three main directions by focusing on privacy. In summary, this survey presents an extensive literature review of the utmost possibilities of eye tracking in VR and their privacy implications.
@@ I Introduction
Over the last decade, virtual, augmented, and mixed reality (VR/AR/MR) communities have benefited from advancements in computer hardware, graphics, and imaging science. To date, some modern head-mounted displays (HMDs) have become available at reasonable prices for everyday use. Furthermore, eye-tracking sensors have become either directly integrated into these HMDs (e.g., HTC Vive Pro Eye [1] or are available as low-cost add-ons (e.g., Pupil Labs [2]). Since eye movements are related to human cognition, visual attention, and perception, the information on where users look can be beneficial to them in various ways, such as providing adaptive support in VR, increasing their engagement, or enhancing the usability of VR applications. At the same time, there is also considerable room for technical improvements, such as enhancing display resolutions, increasing realism, or preventing cybersickness, in which gaze information may be helpful.
It was foreseen in 2017 that VR was approximately 20 years away from achieving resolutions comparable to those of the human eye [3]. With such advancements and a wider range of applications, future HMDs may evolve into mobile devices, similar to today’s mobile phones and tablets. However, as HMDs are situated on the users’ heads and have close proximity to their eyes, they may be perceived as privacy-invasive, as many sensitive attributes about users, such as gender, sexual preferences, and identities, can be extracted using eye movements [4]. From a privacy perspective, this information should be protected, or the user should have control over its release. At the same time, a pleasant virtual experience should be available to all users. Furthermore, eye information can also be utilized for accurate authentication, especially with iris textures, as they are analogous to fingerprints. Therefore, there is an essential need for a privacy-utility trade-off.
As VR and privacy communities have been working to tackle these issues, and with the availability of HMDs to wider communities, especially in recent years, the number of works at the intersection of VR, eye tracking, and privacy has been increasing. However, to date, none of the previous works have focused on comprehensive coverage of eye tracking in VR and its privacy implications, considering the leading venues in the field. In this work, (1) we first cover the research between 2012 and 2022 by discussing how eye movements are extracted and the possibilities with them in immersive VR. Then, (2) we discuss security and privacy implications, including eye-based authentication and privacy-preserving eye tracking. Lastly, (3) we draw and discuss three different directions for future research, especially focusing on privacy.
@@ I-A Paper’s Aim, Structure, and a Navigational Guide
This paper aims to provide researchers and practitioners with a detailed overview of the different aspects of eye tracking in VR. We begin with details on existing hardware, datasets, and algorithms, the latter being more specific to eye-region segmentation and gaze estimation. We then dig deeper into eye-based interaction techniques in VR, which often inherently utilize estimated eye regions and gaze. Furthermore, the overview of works on human cognition, visual attention, and perception in VR aims to teach the reader about the relationship between human eye movements and physiology. The security and privacy aspects discussed in the paper aim to provide methodological details on the relationship between eye movements and user authentication, and how to protect users’ privacy when the focus is not on authentication. Lastly, future directions involve a privacy-aware approach to eye movement-based methods in VR. Overall, the primary objective of this paper is to provide a comprehensive guide to navigating the landscape of eye tracking in VR, encompassing its technical foundations, practical applications, and ethical considerations, while also suggesting directions for further exploration, particularly in the context of privacy.
@@
As we comprehensively survey research conducted in eye tracking, VR, and privacy domains, and due to the interdisciplinary nature of each domain, we provide a navigational guideline on how readers with different backgrounds can benefit from it. We first discuss existing works in the literature in Section II. We then discuss our search methodology, considering the most important academic venues in Section III. These sections are suitable for all readers, particularly those seeking to understand existing surveys and this work’s novelty.
@@
Section IV discusses a range of topics, from computer vision-based approaches for tracking eyes and further processing eye region data to eye-based human-computer interaction (HCI) and understanding human visual attention, cognition, and perception. Section IV-A will benefit most readers interested in computer vision and machine learning. Section IV-B focuses on HCI by utilizing eye movements for various purposes, including real-time interaction in VR, foveated rendering, and addressing technical issues to enhance immersive experiences, which is suitable for readers who work on the technical aspects of HCI and eye tracking. Section IV-C combines eye tracking and VR from a human cognition, attention, and perception perspective, and is suitable for those with cognitive and experimental psychology backgrounds, or those interested in conducting human factors research.
Section V presents works exploring the privacy and security implications of eye-tracked VR, including authentication possibilities, the importance of privacy-preserving methods, and how current literature has addressed privacy issues, which is relevant for both privacy and security researchers in HCI. Lastly, although we primarily discuss the future directions and implications of eye-tracked VR in Section VI with a focus on privacy, this section is relevant for all researchers whose work involves VR and eye tracking, as privacy issues should be carefully handled regardless of the research area. Section VII concludes our paper. In summary, after reading this paper in whole or in part, we anticipate that it will be easier for novice individuals to become knowledgeable about the state-of-the-art, while our paper serves as a concise reference for experienced researchers. Figure 1 provides a holistic visual representation of the concepts discussed in this paper.
Figure 1: Visual representation of eye-tracked VR, considering data capturing, processing, data types, and utility tasks.
@@ II Related work
Several works analyze relevant research for eye tracking, VR/AR, privacy, and security. None systematically and comprehensively reviews eye-tracking in VR setups by covering the major venues, the privacy issues that such works could lead to, and how to mitigate such risks. Previously, [5] analyzed eye tracking in the wild by discussing the advantages, disadvantages, and technical terms to achieve high-quality results from eye-tracking experiments. Furthermore, [6] covered gaze-based interaction and eye tracking in head-worn extended reality (XR). Similarly, [7] discussed the foundations of eye-tracking support for visual analytics systems, applications, and challenges by five critical themes, including privacy protection as an essential issue.
Previous work has studied the challenges in immersive analytics by considering VR, MR, and AR with future directions [8], sector- and domain-specific reviews, such as for sports, highlighting the importance of usability, interaction, and image presentation [9]. Furthermore, [10] examined eye-tracking research in VR, with a pilot user study that focused on experimental setups and common problems, such as vergence-accommodation conflict and eye-tracking calibration, on a small scale. [11] broadly reviewed eye-tracking applications in VR, and similar to [7], they identified privacy and security as essential discussion points. However, their review was not comprehensive and covered privacy as an application with a handful of works. A similar review focused on eye-tracking processing is given in [12], and [13] briefly overviews the privacy aspects. Furthermore, overviews on high-quality eye-tracking data collection [14] and the measurement of data quality in HMD-based eye tracking [15] also exist.
@@
While visual scanning patterns and eye-tracking data provide useful information about users, this information also raises concerns about users’ privacy. [4] discussed that the end-users may not understand the privacy loss and argued that privacy protection should be in place for gaze and pupillometry. [16] highlighted the attributes that can be inferred from gaze data, similar to [4], and discussed the privacy implications and societal impacts. [17] considered multi-user AR’s security and privacy and found that some users were concerned about the powerful abilities of the eye-tracking enabled interfaces and behavioral tracking, such as an HMD understanding that the user is being attracted to someone due to being unable to stop looking at them. This finding partly aligns with the findings of more recent work [18], which indicate that users are more likely to agree to share their eye-tracking data if the data owner is a governmental health agency or if the purpose is research. Furthermore, [19] provided an extensive survey on gaze-based authentication, encouraging further research on privacy-preserving eye tracking like prior works [11, 7].
@@
Despite several recent attempts, none of the previous works have comprehensively and systematically analyzed the research on eye-tracking data analytics in VR, considering the entire processing pipeline of eye-tracking methods. We consider methods from pupil detection and gaze estimation to human visual attention and cognition for processing and mining eye-tracking data. We further explore works that investigate privacy-preserving methods for eye movements in VR setups, providing novel research directions that should be pursued to enhance the practical applications of VR. Considering the number of sensitive attributes that eye-tracking data can reveal, we take privacy as one of the main pillars of our paper. Therefore, our survey, focusing on these aspects, provides a comprehensive overview for the research community.
@@ III Methodology
We first analyzed the recent surveys and papers that study eye tracking in the context of VR, AR, HCI, and privacy [5, 7, 8, 6, 20, 4, 16, 19]. We considered the papers between 2012-2022 since the VR devices and eye trackers have been becoming prevalent lately. We used Google Scholar to query papers from 35 renowned venues including ACM CHI, IEEE VR, ACM ETRA, IEEE ISMAR, ACM IUI, ACM UIST, ACM ICMI, ACM MobileHCI, ACM MM, ACM VRST, IEEE AIVR, AAAI Conference on Artificial Intelligence, IEEE CVPR, IEEE ICCV, ECCV, NeurIPS/NIPS, ACM SUI, IEEE VIS, PoPETS/PETS, USENIX Security, USENIX SOUPS, NDSS, IEEE Symposium on S&P, ACM CCS, ACM Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), Journal of Eye Movement Research (JEMR), IEEE Transactions on Visualization and Computer Graphics (TVCG), ACM Transactions on Graphics (ToG), IEEE S&P Magazine, Journal of Privacy and Confidentiality, IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Dependable and Secure Computing, Computers & Security (C&S), ACM Transactions on Privacy and Security, and ACM Transactions on Cyber-Physical Systems. For the conference proceedings, we considered the main proceedings, excluding companion and workshop proceedings, except for works that present concrete results or are highly related to eye tracking and VR. Proceedings of the ACM ETRA in 2022 were published in the Proceedings of the ACM on Human-Computer Interaction and the ACM on Computer Graphics and Interactive Techniques, and we also included those in our search results under the ETRA category. We cross-checked the query results from the DBLP.
To filter the papers, we used (“eye tracking” AND “virtual reality”) as the main query, where each phrase could appear anywhere in the papers. In addition, we used the following sub-queries to further track privacy- and biometrics-related papers: (“eye tracking” AND “virtual reality” AND “privacy”) and (“eye tracking” AND “virtual reality” AND “biometrics”). We also applied forward and backward tracking by reviewing the citations and references of the queried papers, respectively. The main query yielded 1364 results, whereas two sub-queries led to 215 and 96 results, respectively. Figure 2 shows the resulting distributions of the main query. If a venue above does not appear in Figure 2, it means that we did not find any paper at this venue with our query.
Figure 2: Query responses according to venues for our main query.
@@ IV Eye-tracked Virtual Reality
Eye tracking can be considered one of the key technologies, especially for VR, as it enables the accurate tracking of attention within HMDs. In addition, since eyes do not move completely voluntarily, one can obtain more objective user behaviors compared to self-reported data, such as questionnaires. While one can utilize this rich source of information in various ways, there are different technical steps to achieve this. Firstly, eye regions and gaze directions are extracted primarily using computer vision and machine learning techniques. Later, gaze directions are utilized for interactions, for instance, to support users by predicting their intents and providing them with context-sensitive aid or technical ways, such as foveated rendering. While such cases should work in real-time, collected data from different experimental setups can also be used offline to understand humans and how they behave in various application domains, such as education or medicine, showing that VR can also be considered a research tool to create scenarios that are almost impossible in the wild due to privacy and safety issues or because of practical infeasibility.
Considering all, we organize this section as follows. We first discuss works that utilize computer vision and machine learning to estimate gaze and detect eye regions, as these are the initial steps in data processing, with a significant portion of recent research on these topics employing deep learning-based approaches. We also include eye movement event detection methods in VR, which differ from those in conventional settings. We then provide a detailed overview of eye-based interaction and its various applications, with a particular focus on efficiency, usability, and real-time working capability. As the last step, we cluster the works that conduct offline data analyses, particularly to understand human visual attention, cognition, and perception, and we provide an overarching view of these works. While the offline understanding of human visual attention, cognition, and perception does not have a direct and real-time impact on the interaction experience, understanding these aspects helps to achieve an optimal virtual space and interaction design.